1.Receivables Investment Snider Industries sells on terms of 2/10, net 35. Total sales for the year are $500,000. Thirty percen2.Cost of Trade Credit A large retailer obtains merchandise under the credit terms of 3/10, net 45, but routinely takes 55 days
Answer #1

1)

D13 x fic =+D4-D10 A B с D E 1 a) Days of Sales Outstanding (DSO) 2 31.0 =10*30%+40*70% Days 3 (DSOxSales)/365 $ 42,466 =D1*5

*Please rate thumbs up